# Simulated Bifurcation Quantum Annealing

### 01 Apr 2026 v1

@@ I Introduction
FIG. 1. : Heatmaps of optimality gaps $g$ in the $\alpha,\beta$ plane, demonstrating the sensitivity of the SBQA algorithm. Panels (a)–(f) correspond to different instances; see main text for details. Plotted values of $g$ are averaged over best energies from 10 independent runs for each instance, with $8$ sets of $128$ interacting replicas per run. Red contour lines correspond to a $35\%$ threshold relative to the best value obtained. The results are consistent with interpreting $\beta$ as the inverse temperature, since energies decrease with increasing $\beta$. They suggest selecting $\beta$ from the range $0.5\lesssim\beta\lesssim 1.5$. The impact of $\alpha$ is less pronounced, but still visible, especially for the all-to-all instance in panel (a). We therefore restrict its range to $0.5\lesssim\alpha\lesssim 1.0$. Altogether, this analysis suggests that the SBQA algorithm is stable with respect to the choice of hyperparameters, and no complicated, instance-dependent tuning is required.
@@ II.1 Simulated Bifurcation Quantum Annealing
Our starting point is a classical Hamiltonian describing a system of $N$ particles, interacting via Ising-like interactions, and subject to a time-dependent external potential:
@@
where $a(t)=t/T$ and $T$ is the total time of the evolution. The trajectories of the particles are confined to an $N$-dimensional unit hypercube, which is achieved by imposing that $V=\infty$ if any $q_{i}>1$. Using Hamilton’s equations, we can derive the following dynamical system, called the Simulated Bifurcation Machine (SBM):
@@
Our modification of SBM, which we call the Simulated Bifurcation Quantum Annealing (SBQA), is inspired by the Simulated Quantum Annealing (SQA), also known as Path Integral Monte Carlo [42]. The idea is to couple the previously independent replicas of the system, via a certain time-dependent interaction derived from the Suzuki-Trotter decomposition of the quantum partition function of the transverse-field Ising model, which describes the quantum annealing process. For a composite system of $N$ particles and $R$ replicas, the SBQA Hamiltonian reads:
@@ II.2 Parameter sensitivity and auto-tuning
where $E_{0}$ is the best known energy (in particular the ground state energy, if known) and $E$ is the energy of the returned solution. We plot the results in Fig. I, where each point is averaged over $10$ independent runs, with $8$ sets of $128$ interacting replicas per run. Based on these results, we restrict the range of hyperparameters to $\beta\in[0.5,1.5]$ and $\alpha\in[0.5,1.0]$. Since it is neither obvious how to select the optimal values within these ranges, nor desirable to perform costly, per-instance tuning, we introduce a simple auto-tuning procedure. The total number of processed samples $N_{\rm samples}$ is split into $N_{\rm repetitions}$ with interacting $N_{\rm replicas}$ replicas each. For each repetition, we randomly select $\beta$ and $\alpha$ from the above ranges, and evolve the system. Finally, we return the best solution from all repetitions. This procedure is simple, yet effective, as it allows one to explore a variety of hyperparameter combinations, without incurring significant overhead, since the parallel capabilities of modern GPUs allow all $N_{\rm repetitions}$ repetitions to be executed in parallel.
@@ III.1.1 Large-scale Zephyr instances
In this section, we focus on the Zephyr graphs across four orders of magnitude in size, from $Z_{1}$ with $N=47$ spins, through $Z_{12}$ with $N=4800$ spins (current generation Advantage2 QPU) up to $Z_{150}$ with $N=722400$ spins. Assuming a doubling period of 2 years, a $Z_{150}$ working graph is expected to be within reach of quantum annealers in approximately a decade [30]. For each of the graphs, we generate $20$ instances in the small regime ($Z_{1}-Z_{20}$) and $5$ in the large regime ($Z_{30}-Z_{150}$), with random couplings $J_{ij}$ and fields $h_{i}$ obtained from a QUBO matrix $Q_{ij}$ with entries drawn from a uniform distribution over [1,1][-1,1] (see Ref. [30] for details). The reference energies are obtained by running large scale Simulated Annealing (SA) calculations, operating on timescales typically at least an order of magnitude larger than those required by SBM and SBQA.
@@ III.1.2 Large-scale Quantum Annealing Correction (QAC) problems
FIG. 8. : Average optimality gap $g$ as a function of the number of steps for instances on a cubic lattice with linear dimension (a) $L=10$ and (b) $L=50$. Results are averaged over an ensemble of $10$ random instances for each size and $5$ independent runs per instance; error bars show one standard deviation. Insets in panels (a) and (b) show the measured runtime as a function of the number of steps. Panel (c) shows the difference $\Delta g$ between the optimality gaps obtained by SBM and SBQA as a function of instance size, with negative values favoring SBQA. The advantage of SBQA emerges for $L>20$, but it is less pronounced than in the square-lattice case.
@@ III.2 Instances on current-generation quantum hardware
We start with benchmarks on 3D spin glasses, with couplings distributed according to the standard normal distribution. This is a modification of the instances already studied in the context of combinatorial optimization and quantum annealing, for which both the D-Wave annealer and DTSQA performed very well [22, 6]. The logical topology of the instances is a cubic lattice with periodic boundaries in two directions and an open boundary in the third direction. Such graphs have a simple embedding into the Pegasus architecture, with the $P_{M}$ graph accommodating a lattice of dimensions at most $(M-1)\times(M-1)\times 12$, using chains of fixed length $2$ [3]. We consider cubic lattices of size $L^{3}$ for $L=6,8,10,12$ and $L^{2}\times 12$ for $L=15$. For the reference energies, we used the DTSQA algorithm run with $5\times 10^{5}$ steps. We note that DTSQA requires graph coloring before the computation can be run, and this preprocessing can be performed once for a given instance topology. In our implementation, we use the DSATUR algorithm and exclude the coloring time from the runtime measurements.
In Fig. III.2.1, we compare the SBQA time-to-epsilon values for embedded instances of different sizes against our baseline SBM, DTSQA, and the D-Wave Advantage 4.1 quantum annealer. For each data point, we consider $50$ random instances ($10$ for D-Wave) and $5$ independent runs ($10$ for D-Wave) per instance to estimate the success probability and average runtime (QPU access time for the annealer). As usual, the time-to-epsilon values are minimized over the number of steps of the algorithm, or the annealing time in the case of D-Wave.
@@ Acknowledgments
This project was supported by the National Science Center (NCN), Poland, under Projects: Sonata Bis 10, No. 2020/38/E/ST3/00269 (B.G.). Quantumz.io Sp. z o.o. acknowledges support received from The National Centre for Research and Development (NCBR), Poland, under Project No. POIR.01.01.01-00-0061/22.
@@
Here, we analyze the two components of the time-to-epsilon benchmark for 3D spin glasses, namely the solution quality, measured by the obtained optimality gap $g$, and the runtime of the algorithms. Since the true ground state energies of these instances are unknown, we estimate them by running DTSQA with a very large number of steps, $N_{s}=5\times 10^{5}$, and take the best energy found across $5$ runs as the reference energy $E_{0}$. Crucially, the number of steps used for GS estimation is significantly larger than the maximum number of steps used in the actual benchmarks.

### 07 Apr 2026 v2

@@ I Introduction
FIG. 1. : Heatmaps of optimality gaps $g$ in the $\alpha,\beta$ plane, demonstrating the sensitivity of the SBQA algorithm. Panels (a)–(f) correspond to different instances; see main text for details. Plotted values of $g$ are averaged over best energies from 10 independent runs for each instance, with 88 sets of $128$ interacting replicas per run. Red contour lines correspond to a $35\%$ threshold relative to the best value obtained. The results are consistent with interpreting $\beta$ as the inverse temperature, since energies decrease with increasing $\beta$. They suggest selecting $\beta$ from the range $0.5\lesssim\beta\lesssim 1.5$. The impact of $\alpha$ is less pronounced, but still visible, especially for the all-to-all instance in panel (a). We therefore restrict its range to $0.5\lesssim\alpha\lesssim 1.0$. Altogether, this analysis suggests that the SBQA algorithm is stable with respect to the choice of hyperparameters, and no complicated, instance-dependent tuning is required.
@@ II.1 Simulated Bifurcation Quantum Annealing
Our starting point is a classical Hamiltonian describing a system of NN particles, interacting via Ising-like interactions, and subject to a time-dependent external potential:
@@
where $a(t)=t/T$ and $T$ is the total time of the evolution. The trajectories of the particles are confined to an NN-dimensional unit hypercube, which is achieved by imposing that $V=\infty$ if any $q_{i}>1$. Using Hamilton’s equations, we can derive the following dynamical system, called the Simulated Bifurcation Machine (SBM):
@@
Our modification of SBM, which we call the Simulated Bifurcation Quantum Annealing (SBQA), is inspired by the Simulated Quantum Annealing (SQA), also known as Path Integral Monte Carlo [42]. The idea is to couple the previously independent replicas of the system, via a certain time-dependent interaction derived from the Suzuki-Trotter decomposition of the quantum partition function of the transverse-field Ising model, which describes the quantum annealing process. For a composite system of NN particles and $R$ replicas, the SBQA Hamiltonian reads:
@@ II.2 Parameter sensitivity and auto-tuning
where $E_{0}$ is the best known energy (in particular the ground state energy, if known) and $E$ is the energy of the returned solution. We plot the results in Fig. I, where each point is averaged over $10$ independent runs, with 88 sets of $128$ interacting replicas per run. Based on these results, we restrict the range of hyperparameters to $\beta\in[0.5,1.5]$ and $\alpha\in[0.5,1.0]$. Since it is neither obvious how to select the optimal values within these ranges, nor desirable to perform costly, per-instance tuning, we introduce a simple auto-tuning procedure. The total number of processed samples $N_{\rm samples}$ is split into $N_{\rm repetitions}$ with interacting $N_{\rm replicas}$ replicas each. For each repetition, we randomly select $\beta$ and $\alpha$ from the above ranges, and evolve the system. Finally, we return the best solution from all repetitions. This procedure is simple, yet effective, as it allows one to explore a variety of hyperparameter combinations, without incurring significant overhead, since the parallel capabilities of modern GPUs allow all $N_{\rm repetitions}$ repetitions to be executed in parallel.
@@ III.1.1 Large-scale Zephyr instances
In this section, we focus on the Zephyr graphs across four orders of magnitude in size, from $Z_{1}$ with $N=47$ spins, through $Z_{12}$ with $N=4800$ spins (current generation Advantage2 QPU) up to $Z_{150}$ with $N=722400$ spins. Assuming a doubling period of 2 years, a $Z_{150}$ working graph is expected to be within reach of quantum annealers in approximately a decade [30]. For each of the graphs, we generate 2020 instances in the small regime ($Z_{1}-Z_{20}$) and 55 in the large regime ($Z_{30}-Z_{150}$), with random couplings $J_{ij}$ and fields $h_{i}$ obtained from a QUBO matrix $Q_{ij}$ with entries drawn from a uniform distribution over [1,1][-1,1] (see Ref. [30] for details). The reference energies are obtained by running large scale Simulated Annealing (SA) calculations, operating on timescales typically at least an order of magnitude larger than those required by SBM and SBQA.
@@ III.1.2 Large-scale Quantum Annealing Correction (QAC) problems
FIG. 8. : Average optimality gap $g$ as a function of the number of steps for instances on a cubic lattice with linear dimension (a) $L=10$ and (b) $L=50$. Results are averaged over an ensemble of $10$ random instances for each size and 55 independent runs per instance; error bars show one standard deviation. Insets in panels (a) and (b) show the measured runtime as a function of the number of steps. Panel (c) shows the difference $\Delta g$ between the optimality gaps obtained by SBM and SBQA as a function of instance size, with negative values favoring SBQA. The advantage of SBQA emerges for $L>20$, but it is less pronounced than in the square-lattice case.
@@ III.2 Instances on current-generation quantum hardware
We start with benchmarks on 3D spin glasses, with couplings distributed according to the standard normal distribution. This is a modification of the instances already studied in the context of combinatorial optimization and quantum annealing, for which both the D-Wave annealer and DTSQA performed very well [22, 6]. The logical topology of the instances is a cubic lattice with periodic boundaries in two directions and an open boundary in the third direction. Such graphs have a simple embedding into the Pegasus architecture, with the $P_{M}$ graph accommodating a lattice of dimensions at most $(M-1)\times(M-1)\times 12$, using chains of fixed length 22 [3]. We consider cubic lattices of size $L^{3}$ for $L=6,8,10,12$ and $L^{2}\times 12$ for $L=15$. For the reference energies, we used the DTSQA algorithm run with $5\times 10^{5}$ steps. We note that DTSQA requires graph coloring before the computation can be run, and this preprocessing can be performed once for a given instance topology. In our implementation, we use the DSATUR algorithm and exclude the coloring time from the runtime measurements.
In Fig. III.2.1, we compare the SBQA time-to-epsilon values for embedded instances of different sizes against our baseline SBM, DTSQA, and the D-Wave Advantage 4.1 quantum annealer. For each data point, we consider $50$ random instances ($10$ for D-Wave) and 55 independent runs ($10$ for D-Wave) per instance to estimate the success probability and average runtime (QPU access time for the annealer). As usual, the time-to-epsilon values are minimized over the number of steps of the algorithm, or the annealing time in the case of D-Wave.
@@ Acknowledgments
This project was supported by the National Science Center (NCN), Poland, under Projects: Sonata Bis 10, No. 2020/38/E/ST3/00269 (B.G.) and Sonata Bis 15, No. 2025/58/E/ST6/00422 (Ł.P.). Quantumz.io Sp. z $o.$ acknowledges support received from The National Centre for Research and Development (NCBR), Poland, under Project No. POIR.01.01.01-00-0061/22.
@@
Here, we analyze the two components of the time-to-epsilon benchmark for 3D spin glasses, namely the solution quality, measured by the obtained optimality gap $g$, and the runtime of the algorithms. Since the true ground state energies of these instances are unknown, we estimate them by running DTSQA with a very large number of steps, $N_{s}=5\times 10^{5}$, and take the best energy found across 55 runs as the reference energy $E_{0}$. Crucially, the number of steps used for GS estimation is significantly larger than the maximum number of steps used in the actual benchmarks.